OpenStudy (anonymous):

what is 16/4 as a whole number

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \sf 4 \times 4=16}\)

OpenStudy (anonymous):

well all you have to do is divide 4 into 16........

OpenStudy (anonymous):

so it would be 4

OpenStudy (skullpatrol):

@desiree1996 Are you still there?

OpenStudy (skullpatrol):

$$\Huge a \div b = \frac{a}{b}$$

OpenStudy (skullpatrol):

$$\Huge 16\div 4 = \frac{16}{4}$$

OpenStudy (anonymous):

Ok thanks for the help

OpenStudy (skullpatrol):